OpenStudy (anonymous):

write the equation in its equivalent logarithmic form 9^3=729

OpenStudy (lgbasallote):

\[\huge a^b = c \Longleftrightarrow \log_a c = b\] so what's your a here?

OpenStudy (anonymous):

dont know

OpenStudy (lgbasallote):

well do you agree that \[9^3 = 729\] is in the form \[a^ b = c\]

OpenStudy (anonymous):

ok

OpenStudy (lgbasallote):

it was a question...

OpenStudy (anonymous):

yes sir

OpenStudy (lgbasallote):

lol

OpenStudy (anonymous):

:(

OpenStudy (lgbasallote):

read my 2nd post again

OpenStudy (anonymous):

:)